OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (6, 1), (5, 4) need answer and explanation.

OpenStudy (amistre64):

x moves from 6 to 5; its a movement of -1 y moves for 1 to 4; its a movement of +3 slope = y/x = 3/-1; which can be written as -3/1 as well .................................................................................... ax +by = c ; where slope = -a/b a = -(-3) ; and b=1 which gives us: 3x +y = (some#) ; to find the number insert either point; (5,4) perhaps 3(5) + (4) = 19 to wit; 3x +y = 19